OpenStudy (anonymous):

The weight of a bag of quarters and a bag of dimes is 510.3 g. A quarter weighs 5.67 grams and a dime weighs 2.268 grams. There are 100 dimes. Find the number of quarters.

OpenStudy (sleepyjess):

Ok... How do you think you would start this with the info given?

OpenStudy (anonymous):

By adding all the grams together then subtract that total number by 100

OpenStudy (sleepyjess):

quarters + dimes = 510.3g q(5.67) + 100(2.268) = 510.3

OpenStudy (sleepyjess):

Now what? :-)

OpenStudy (anonymous):

im confused with this probelm

OpenStudy (sleepyjess):

q(5.67) + 226.8 = 510.3 q(5.67) + 226.8 - 226.8 = 510.3 - 226.8 q(5.67) = 283.5 Divide both sides by 5.67 to get the number of quarters on its' own

OpenStudy (anonymous):

50

OpenStudy (sleepyjess):

You have the (number of quarters x their weight) + (number of dimes x their weight) = 510.3g You are given the weight of the quarters, the weight of the dimes and the number of dimes, so you just substitute those into the equation. Yes, that is the correct answer

OpenStudy (sleepyjess):

If you can just put it into terms that you understand first, it makes it a lot easier than it seems at first :-)

OpenStudy (anonymous):

thanks for the advice